A bookstore had 80 copies of a magazine. Yesterday, it sold 2/5 of them. Today, it sold 3/4 of what remained. How many copies does the bookstore have left?
Answer by nerdybill(7384)

A bookstore had 80 copies of a magazine. Yesterday, it sold 2/5 of them. Today, it sold 3/4 of what remained. How many copies does the bookstore have left?
Started with
80
.
Yesterday:
Sold: (2/5)(80) = (2)16 = 32
Remaining: 80-32= 48
.
Today:
Sold: (3/4)(48) = (3)12 = 12
Left over: 48-12 = 36 magazines